# Groups Acting Acylindrically on Trees

### v5

@@ 4.3. A Geometric Condition for Acylindrical Arboreality
Let $M$ be a closed and orientable hyperbolic $3$-manifold with fundamental group $G$. That (1) implies (2) is a standard result for hyperbolic groups, see [Linton22] for example, but we include a brief proof for completeness. Assume that $G$ admits a non-elementary quasi-convex splitting of some quasi-convex subgroup $H$, with corresponding Bass–Serre tree $T$, and recall that by Theorem 4.7 $G$ is hyperbolic relative to the trivial subgroup. Then $H$ is by definition relatively quasi-convex with respect to the trivial subgroup, and so by Theorem 4.2 there exists a natural number $n_{H}$ such that the intersection of at least $n_{H}$ distinct conjugates of $H$ must be a finite or peripheral subgroup. However, all of our peripheral subgroups are trivial in this case, so such an intersection must be finite, and by Lemma 4.10 and the fact that $G$ admits a non-elementary and thus non trivial quasi-convex splitting and is therefore torsion free we must have that $G$ is torsion free. Thus the intersection of at least $n_{H}$ essentially distinct conjugates of $H$ is trivial. Now let $p$ be a path in $T$ of length at least $n_{H}$. The stabiliser for this path will be the intersection of the stabilisers of its constituent edges, which are essentially distinct conjugates of $H$ in $G$, and so $\operatorname{PStab}_{G}(p)$ is trivial, and the action of $G$ on $T$ is therefore $(n_{H},1)$-acylindrical. This action was non-elementary by assumption, and so it follows that $G$ is acylindrically arboreal, and (1) implies (2).
Now assume that (2) holds. By Theorem 1.6, $M$ contains an embedded $2$-sided incompressible closed subsurface $\Sigma$ that is not isotopic to any boundary component of $M$, and such that the image of the natural inclusion $\pi_{1}(\Sigma)\hookrightarrow\pi_{1}(M)$ is geometrically finite. Therefore, by Theorem 4.11, $\pi_{1}(\Sigma)$ must be relatively quasi-convex with respect to the boundary subgroups of $G$. The manifold $M$ has empty boundary, so $G$ is hyperbolic and $\pi_{1}(S)$ is a quasi-convex subgroup of $G$. It follows that the cut of $G$ along $\Sigma$ will induce a quasi-convex splitting of $G$. This splitting will be non-trivial by construction of $\Sigma$, and will be acylindrical by the proof that (1) implies (2). The group $G$ is not virtually cyclic as it contains a non-trivial surface subgroup, and so by Theorem 2.3 the cut of $M$ along $\Sigma$ induces a non-trivial splitting, which will then be a non-trivial quasi-convex splitting as required. It follows that (2) implies (1), and so (1) and (2) are equivalent as claimed.
@@
Let $e$ be any edge of $(\Gamma,\mathfrak{G})$ such that the splitting $(\Gamma^{\prime},\mathfrak{G}^{\prime})$ of $G$ over $G_{e}$ is non-trivial, where $\Gamma^{\prime}$ and $\mathfrak{G}^{\prime}$ are obtained from $\Gamma$ by collapsing all edges that are not $e$ and updating the vertex groups accordingly. Then $\Gamma^{\prime}$ has exactly one edge (labelled $e$) and either one or two vertices, so by Theorem 4.6 there exists at least one compact $2$-sided incompressible surface $\Sigma$ in $M$ that is not boundary parallel such that $\pi_{1}(\Sigma)\leq G$ is contained in some conjugate $G_{e}^{\prime}$ of $G_{e}$.
The subgroup $\pi_{1}(\Sigma)$ is the fundamental group of a compact surface so is finitely generated, and by assumption and by Theorem 1.6, $\pi_{1}(\Sigma)$ is not geometrically finite so it must be a virtual surface fibre subgroup by Theorem 4.12 (subgroup tameness). Thus, by Lemma 4.15, $G_{e}^{\prime}$ is a virtual surface fibre subgroup of $G$, and so $G_{e}$ is conjugate to some virtual surface fibre subgroup of $G$ and is therefore is a virtual surface fibre subgroup itself.
@@
with the last congruence following from the fact that $G_{e}$ has index two in both $\pi_{1}(M_{1})$ and $\pi_{2}(M)$. The group $D_{\infty}$ is virtually cyclic so must act on $\overline{G\cdot e}\subseteq T$ with a fixed point or fixed line, and thus the action of $G$ on $T$ will have a fixed point or fixed line.
@@
    By a result of Floyd and Hatcher [FLOYD1982263, Theorem 1.1] all closed incompressible embedded surfaces in $M$ are isotopic to the torus boundary component. The fundamental group $\pi_{1}(M)$ is therefore not acylindrically arboreal by Theorem 1.6.

### v6

@@ 4.3. A Geometric Condition for Acylindrical Arboreality
Let $M$ be a closed and orientable hyperbolic $3$-manifold with fundamental group $G$. That (1) implies (2) is a standard result for hyperbolic groups, see [undefy] for example, but we include a brief proof for completeness. Assume that $G$ admits a non-elementary quasi-convex splitting over some quasi-convex subgroup $H$, with corresponding Bass–Serre tree $T$, and recall that by Theorem 4.9 $G$ is hyperbolic relative to the trivial subgroup. Then $H$ is by definition relatively quasi-convex with respect to the trivial subgroup, and so by Theorem 4.3 there exists a natural number $n_{H}$ such that the intersection of at least $n_{H}$ distinct conjugates of $H$ must be a finite or peripheral subgroup. However, all of our peripheral subgroups are trivial in this case, so such an intersection must be finite, and by Lemma 4.12 and the fact that $G$ admits a non-elementary and thus non trivial quasi-convex splitting and is therefore torsion free we must have that $G$ is torsion free. Thus the intersection of at least $n_{H}$ essentially distinct conjugates of $H$ is trivial. Now let $p$ be a path in $T$ of length at least $n_{H}$. The stabiliser for this path will be the intersection of the stabilisers of its constituent edges, which are essentially distinct conjugates of $H$ in $G$, and so $\operatorname{PStab}_{G}(p)$ is trivial, and the action of $G$ on $T$ is therefore $(n_{H},1)$-acylindrical. This action was non-elementary by assumption, and so it follows that $G$ is acylindrically arboreal, and (1) implies (2).
Now assume that (2) holds. By Theorem 1.6, $M$ contains an embedded $2$-sided incompressible closed subsurface $\Sigma$ that is not isotopic to any boundary component of $M$, and such that the image of the natural inclusion $\pi_{1}(\Sigma)\hookrightarrow\pi_{1}(M)$ is geometrically finite. Therefore, by Theorem 4.13, $\pi_{1}(\Sigma)$ must be relatively quasi-convex with respect to the boundary subgroups of $G$. The manifold $M$ has empty boundary, so $G$ is hyperbolic and $\pi_{1}(\Sigma)$ is a quasi-convex subgroup of $G$. It follows that the cut of $G$ along $\Sigma$ will induce a quasi-convex splitting of $G$. This splitting will be non-trivial by construction of $\Sigma$, and will be acylindrical by the proof that (1) implies (2). The group $G$ is not virtually cyclic as it contains a non-trivial surface subgroup, and so by Theorem 2.3 the cut of $M$ along $\Sigma$ induces a non-trivial splitting, which will then be a non-trivial quasi-convex splitting as required. It follows that (2) implies (1), and so (1) and (2) are equivalent as claimed.
@@
Let $e$ be any edge of $(\Gamma,\mathfrak{G})$ such that the splitting $(\Gamma^{\prime},\mathfrak{G}^{\prime})$ of $G$ over $G_{e}$ constructed in the proof of Lemma 2.20 is non-trivial. Then $\Gamma^{\prime}$ has exactly one edge (labelled $e$) and either one or two vertices, so by Theorem 4.8 there exists at least one compact $2$-sided incompressible surface $\Sigma$ in $M$ that is not boundary parallel such that $\pi_{1}(\Sigma)\leq G$ is contained in some conjugate $G_{e}^{\prime}$ of $G_{e}$.
The subgroup $\pi_{1}(\Sigma)$ is the fundamental group of a compact surface so is finitely generated. Furthermore, by assumption and by Theorem 1.6, $\pi_{1}(\Sigma)$ is not geometrically finite so therefore must be a virtual surface fibre subgroup by Theorem 4.14 (subgroup tameness). Thus, by Lemma 4.17, $G_{e}^{\prime}$ is either a virtual surface fibre subgroup of $G$ or has finite index in $G$. The latter case cannot occur as $G_{e}$ is an edge group in a non-trivial splitting and so is not finite index and hence not conjugate to a finite index subgroup, and so $G_{e}^{\prime}$ is in fact a virtual surface fibre subgroup of $G$. Therefore $G_{e}$ is conjugate to some virtual surface fibre subgroup of $G$ and is therefore is a virtual surface fibre subgroup itself.
@@
with the last congruence following from the fact that $G_{e}$ has index two in both $\pi_{1}(M_{1})$ and $\pi_{2}(M)$ by Theorem 4.16. The group $D_{\infty}$ is virtually cyclic so must act on $\overline{G\cdot e}\subseteq T$ with a fixed point or fixed line, and thus the action of $G$ on $T$ will have a fixed point or fixed line.
@@
    By a result of Floyd and Hatcher [undefj, Theorem 1.1] all closed incompressible embedded surfaces in $M$ are isotopic to the torus boundary component. The fundamental group $\pi_{1}(M)$ is therefore not acylindrically arboreal by Theorem 1.6.